Write each subtraction expression as an addition expression. Then add.
1. 18-4 _______________________
2. 6-12 _______________________
3. -7 - 13 ______________________
4. 12-(-2) ______________________
5. m-m ___________________________
6. -z - (-z) _____________________


Thanks So much!
~Bella

Answer by stanbon(57413)

Write each subtraction expression as an addition expression. Then add.
1. 18-4 = 18 + (-4) = 14 _______________________
2. 6-12 = 6 + (-12) = -6 _______________________
3. -7 - 13 = -7+(-13) = -20______________________
4. 12-(-2)= 12+2 = 14 ______________________
5. m-m = m + (-m) = 0___________________________
6. -z - (-z) = -z + z = 0
========================
Cheers,
Stan H.